# existence1.bib

### v1

@@ Abstract
We provide new sufficient conditions for the finiteness of the optimal value and existence of solutions to a general problem of minimizing a proper closed function over a nonempty closed set. The conditions require an asymptotically bounded decay of a function, a relaxation of $p$-supercoercivity, and a certain relation for the asymptotic cone of the constraint set and the asymptotic function of the objective function. Our analysis combines these conditions with a regularization technique. We refine the notion of retractive directions of a set, extend its definition to functions, and establish some basic relations for such directions for both sets and functions. Using these tools, we provide existence of solutions results that generalize many of the results in the literature for both non-convex and convex problems.
@@ 1 Introduction
Our interest is in investigating sufficient conditions for the existence of solutions to general non-convex minimization problems. The existence of solutions has been extensively studied starting with seminal work [Frank_Wolfe_Theorem_56] showing that a quadratic function, which is bounded from below, attains its minimum on a polyhedral set. The result has been extended to the problem with a quasi-convex objective function in [Luo_Zhang_1999]. The work in [Belousov_Klatte_2002] has established that a convex polynomial attains its solution on a region described by finitely many convex polynomial inequalities, which in turn generalized the result established in [Terlaky-LP] for convex quadratic functions.
@@
The inherent difficulty in establishing the existence of solutions is due to directions in unbounded constraint set along which the function may decrease. A unifying framework to address the problem of unboundedness in both functions and sets is proposed in [Auslender_1997, auslender_teboulle_2002], relying on concepts such as the asymptotic cone and asymptotic function to show existence and stability results for general classes of optimization problems. Subsequently, in [bertsekas_tseng_2006], these notions were extended to introduce retractive directions and prove existence of solutions via a nonempty level-set intersection approach. Later on, the work in [ozdaglar_tseng_2006] has developed the existence results for problems where the constraint sets are given by functional inequalities. More recently, a solution existence result has been provided in [Hieu_2021] for a general polynomial objective and a closed constrained set under a certain regularity condition.
In this paper, we generalize the aforementioned results by imposing some conditions on the objective function and the constraint set which are weaker than those in the existing literature. In particular, our existence of solution results extend the sufficient conditions of Theorem 3.4.1 in [auslender_teboulle_2002] for an unconstrained problem to a constrained problem. While the work in [bertsekas_tseng_2006] and [ozdaglar_tseng_2006] have been aimed at the same type of extension, our results are more general and, in particular, they recover Propositions 12 and 13 in [bertsekas_tseng_2006], while extending the class of problems to which Proposition 3.1 in [ozdaglar_tseng_2006] can be applied.
Our development is based on three main concepts, as follows. (1) The class of functions that have asymptotically bounded decay, which is inspired by the super-coercivity [Cambini-Carosi-Coercive, Lara_2022]. This class of functions is wide and includes, for example, polynomials, convex functions, and functions with Lipschitz continuous $p$-derivative for some $p\geq 0$. (2) The asymptotic cone of a set [auslender_teboulle_2002] and the asymptotic cone of a function, which extends such a notion for a proper convex function, as introduced in Definition 2.5.2 of [auslender_teboulle_2002], to any proper function. (3) The cone of retractive directions of a set and a function. The notion of retractive direction of a set builds on Definition 2.3.1 in [auslender_teboulle_2002], and it is slightly more general than a related Definition 1 in [bertsekas_tseng_2006]. The cone of retractive directions of a function is a new concept to the best of our knowledge. We explore the basic properties of the retractive directions of a set and a function.
As a first result, we provide a necessary condition for the finiteness of an optimal value for a constrained problem. Then, using the aforementioned main concepts, we establish our main results for the existence of solutions for problems where the constraint set is non-algebraic. Then, we refine the results for convex problems for cases when the constraint set is non-algebraic and algebraic. Finally, we extend the main results to non-convex problems where the constraint set is specified by functional constraints.
@@
This paper is organized as follows: Section 2 provides the necessary background required for the subsequent development. Section 3 introduces the notion of asymptotically bounded decay of a function, the retractive directions of a set and a function, and investigates their properties. Section 4 presents our main results and their proofs for the problems where the constraint set is generic and not specified via inequalities. Section 5 refines our main results for convex problems and compares them with the existing results in the literature. Finally, Section 6 further elaborates our main results for the case when the constraint set is specified by functional inequalities. The results are also compared with the closely related results reported in [bertsekas_tseng_2006, ozdaglar_tseng_2006].
@@ 2 Notation and Terminology
2.1 Basic Definitions
We consider the functions that take values in the set {$\mathbb{R}\cup\{+\infty\}$. For a function $f$, we use ${\rm dom}f$ to denote its effective domain, i.e., ${\rm dom}f=\{x\in\mathbb{R}^{n}\mid f(x)<+\infty\}$. The epigraph of a function $f$ is denoted by ${\rm epi}f$, i.e., ${\rm epi}f=\{(x,c)\in\mathbb{R}^{n}\times\mathbb{R}\mid f(x)\leq c\}$. The following standard definitions are used repeatedly.
Definition 2.1.1.
For any $\gamma\in\mathbb{R}$, the lower-level set of a function $f:\mathbb{R}^{n}\rightarrow\mathbb{R}\cup\{+\infty\}~{}$is given by
$L_{\gamma}(f)=\{x\in\mathbb{R}^{n}\mid f(x)\leq\gamma\}.$
Our focus is on proper and closed functions, which are defined as follows.
Definition 2.1.2.
The function $f:\mathbb{R}^{n}\rightarrow\mathbb{R}\cup\{+\infty\}~{}$is said to be proper if its epigraph is a nonempty set.
Definition 2.1.3.
The function $f$ is said to be closed if its epigraph ${\rm epi}f$ is a closed set.
The $p$-supercoercivity has been introduced in [Cambini-Carosi-Coercive, Lara_2022], which states that: a proper function $f$ is $p$-supercoercive, with $p\geq 1$, if
$\liminf_{\|x\|\rightarrow\infty}\frac{f(x)}{\|x\|^{p}}>0.$
When the preceding relation holds with $p=0$, the function is coercive, a property that has been widely used.
@@ 2.2 Asymptotic Cones
$x$$y$
Figure 2.1: The set $X=\{(x,y)\in\mathbb{R}^{2}\mid x^{2}\leq|y|\}$ and its asymptotic cone $X_{\infty}=\{(0,y)\mid y\in\mathbb{R}\}$.
$x$$y$
Figure 2.2: The set $X=\{(x,y)\in\mathbb{R}^{2}\mid x^{2}\geq y\}$ and its asymptotic cone $X_{\infty}=\mathbb{R}^{2}$.
We turn our attention to unbounded sets whose behavior at infinity is captured by their asymptotic cones. We provide the definitions of a sequence that converges in a direction and the asymptotic cone of a set, as given in [auslender_teboulle_2002].
Definition 2.2.1.
A sequence $\{x_{k}\}\subset\mathbb{R}^{n}$ is said to converge in the direction $d\in\mathbb{R}^{n}$ if there exists a scalar sequence $\{t_{k}\}\subset\mathbb{R}$ with $t_{k}\rightarrow+\infty$ such that
$\lim_{k\rightarrow\infty}\frac{x_{k}}{t_{k}}=d.$
Definition 2.2.2.
@@
For a nonempty set $X$, the set $X_{\infty}$ is a closed cone by Proposition 2.1.1 in [auslender_teboulle_2002]. Asymptotic cones are illustrated in Figure 2.1 and Figure 2.2. We conclude this section with a result given by Proposition 2.1.9 in [auslender_teboulle_2002].
Proposition 2.2.1.
@@
$(\cap_{i\in\mathcal{I}}C_{i})_{\infty}\subseteq\cap_{i\in\mathcal{I}}(C_{i})_{\infty}\qquad\hbox{whenever $\cap_{i\in\mathcal{I}}C_{i}$ nonempty}.$
@@ 2.3 Asymptotic Functions
Consider the concept of an asymptotic cone applied to the epigraph of a proper function. Doing so allows us to characterize the related notion of asymptotic functions. The formal definition is as follows, according to [auslender_teboulle_2002].
Definition 2.3.1.
@@
A useful analytic representation of an asymptotic function $f_{\infty}$ was originally obtained in [Auslender_1997] and, also, given in Theorem 2.5.1 of [auslender_teboulle_2002].
Theorem 2.3.1.
@@
$f_{\infty}(d)=\liminf_{\begin{subarray}{c}d^{\prime}\rightarrow d\\ t\rightarrow+\infty\end{subarray}}\frac{f(td^{\prime})}{t}$ (1)
or, equivalently,
@@
An asymptotic function has some basic properties inherent from its definition, namely that $f_{\infty}$ is closed and positively homogeneous function since its epigraph is the closed cone $({\rm epi}f)_{\infty}$. Further, the value $f_{\infty}(0)$ is either finite or $f_{\infty}(0)=-\infty$. If $f_{\infty}(0)$ is finite, then it must be that $f_{\infty}(0)=0$ by the positive homogeneity property. As a consequence, for a proper function $f$ we have that $0\in\{d\mid f_{\infty}(d)\leq 0\}$, implying that
$\{d\mid f_{\infty}(d)\leq 0\}\neq\emptyset.$
The directions $d$ such that $f_{\infty}(d)\leq 0$ will be particularly important in our subsequent development. To this end, we will term them as asymptotic directions of a function, and use these directions to define the asymptotic cone of a function, as follows.
Definition 2.3.2.
@@
An asymptotic direction of a function has been given in Definition 3.1.2 of [auslender_teboulle_2002], while the asymptotic cone of a function has been defined for a proper convex function in Definition 2.5.2 of [auslender_teboulle_2002]. However, we adopt the same definition for an arbitrary proper function.
We next provide a key result for the asymptotic cones of lower-level sets of a proper function, which we will use later on. The result can be found in [auslender_teboulle_2002].
Proposition 2.3.1 (Proposition 2.5.3, [auslender_teboulle_2002]).
For a proper function $f$ and any $\alpha\in\mathbb{R}$ such that $L_{\alpha}(f)\neq\emptyset$, one has $\left(L_{\alpha}(f)\right)_{\infty}\subseteq L_{0}(f_{\infty})$ i.e.,
@@
We conclude this section with an existence of solutions result that will also be used in the sequel.
Proposition 2.3.2.
Let $X\subseteq\mathbb{R}^{n}$ be a nonempty closed set, and let $f:\mathbb{R}^{n}\rightarrow\mathbb{R}\cup\{+\infty\}~{}$be a proper closed function with $X\cap{\rm dom}f\neq\emptyset$. If $f$ is coercive over $X$, i.e.,
@@
The result follows by applying Theorem 2.14 of [beck-book] to the function $f+\delta_{X}$, where $\delta_{X}$ is the characteristic function of the set $X$. ∎∎
3 Problem Setup and Basic Concepts
In this section, we introduce the non-convex minimization problem of interest along with new concepts that we will use in the development of our main solution existence results. The problem we consider is
$\inf_{x\in X}f(x),$ ($P$)
where $X\subseteq\mathbb{R}^{n}$ is a nonempty closed set and $f$ is a proper closed function. We let $f^{*}$ denote the optimal value of the problem, and $X^{*}$ denote the set of optimal solutions.
In what follows, we consider a set $X$ that is unbounded. In Subsection 3.1, we introduce a notion of asymptotically bounded decay of a function by bounding the asymptotic behavior of the ratio $f(x)/g(x)$ for an arbitrary function $g$ on $X$. In Subsections 3.2 and 3.3, we introduce the notions of retractive directions for sets and functions, respectively.
@@ 3.1 Asymptotically Bounded Decay
In this section, we introduce a condition on the function $f$ that generalizes the coercivity property. The formal definition is as follows.
Definition 3.1.1.
@@
Note that if were to choose $g(x)=\|x\|^{p}$ for $p\geq 1$, and
$\liminf_{\|x\|\rightarrow\infty\atop x\in X}\frac{f(x)}{\|x\|^{p}}>0,$
relation (3) implies that $f$ is $p$-supercoercive on $X$ and, thus, coercive on $X$. The condition in (3) prohibits the function $f$ from approaching $\infty$- ∞, along the points in the set $X$, faster than the function $g$. The class of functions that have this property is wide. Below we provide several examples.
@@
Hence, any proper function with a finite minimum on a set $X$ satisfies Definition 3.1.1 with $g(x)=\|x\|$. ∎
Next, we show that any proper convex function exhibits asymptotically bounded decay with respect to $g(x)=\|x\|$ due to the special linear underestimation property of a convex function.
@@
Let $f$ be a proper convex function such that $X\cap{\rm dom}f\neq\emptyset$, and let $x_{0}$ be a point in the relative interior of ${\rm dom}f$. Then, by Theorem 23.4 of [rockafellar-1970a], the subdifferential set $\partial f(x_{0})$ is nonempty. Thus, by the convexity of $f$ we have for a subgradient $s_{0}$ of $f$ at the point $x_{0}$ and for all $x\in X$,
@@
Our next example shows that if a function $f$ satisfies $f_{\infty}(d)\geq 0$ for all nonzero $d\in\mathbb{R}^{n}$, then $f$ exhibits asymptotically bounded decay with respect to $g(x)=\|x\|$.
@@
$\liminf_{\|x\|\to\infty\atop x\in X}\frac{f(x)}{\|x\|}=\liminf_{\begin{subarray}{c}d=x\|x\|^{-1}\\ \|x\|\to\infty\atop x\in X\end{subarray}}\frac{f(\|x\|d)}{\|x\|}\geq\liminf_{\begin{subarray}{c}d^{\prime}\to d\\ \|x\|\to\infty\end{subarray}}\frac{f(\|x\|d^{\prime})}{\|x\|}.$
We further have
$\liminf_{\begin{subarray}{c}d^{\prime}\to d\\ \|x\|\to\infty\end{subarray}}\frac{f(\|x\|d^{\prime})}{\|x\|}\geq\liminf_{\begin{subarray}{c}d^{\prime}\to d\\ t\to\infty\end{subarray}}\frac{f(td^{\prime})}{t}=f_{\infty}(d)\geq 0.$
Hence,
$\liminf_{\|x\|\to\infty}\frac{f(x)}{\|x\|}\geq 0,$
and $f$ exhibits asymptotically bounded decay with respect to $g(x)=\|x\|$. ∎
Lastly, we show that a function that is $p$-times differentiable with a Lipschitz continuous $p$th differentials exhibits asymptotically bounded decay with respect to $g(x)=\|x\|^{p+1}$ on $X$.
@@
where $D^{p}f(x)$ denotes the $p$th derivative of $f$ at a point $x$. Then, by Equation (1.5) of  [nesterov2021] we have that
@@
where we used the fact that $\lim_{\|x\|\to\infty}\Phi_{x_{0},p}(x)/\|x\|^{p+1}=0$. Hence, $f$ exhibits asymptotically bounded decay with respect to $g(x)=\|x\|^{p+1}$ on $X$. ∎
Note that multivariate polynomials are a special class of functions that fall under Example 3.4. In particular, a multivariate polynomial of order $m$, with $m\geq 1$, has a constant $m$th order derivative so it is bounded by some constant $B$. Thus, the $(m-1)$st derivative is Lipshitz continuous with the constant $B$. According to Example 3.4, a multivariate polynomial exhibits asymptotically bounded decay with respect to the function $g(x)=\|x\|^{m}$.
3.2 Retractive Directions of Sets

### v3

# On existence of solutions to nonconvex minimization problems

@@ Abstract
We provide a unified framework for a systematic analysis of the existence of solutions to general nonconvex problems, relying on asymptotic and retractive cones for functions and sets. Using this framework we develop new necessary and sufficient conditions for the existence of solutions to a general problem of minimizing a proper closed function over a closed, possibly unbounded, set. Towards the result, we introduce cones of retractive directions for a set and a function, establishing some basic properties for them. We also investigate the relationships between the cone of retractive directions of a function and the cone of level sets of the function. Using the cones of retractive directions we provide necessary and sufficient conditions for the existence of solutions that require an asymptotically bounded decay of a function, and a relation between the cones of retractive directions of the constraint set and the asymptotic cone of the objective function. Finally we refine these conditions for more structured problems.
keywords:
Existence of Solutions, Nonconvex Problems, Optimization, Asymptotic Cone, Asymptotic Function
@@ 1 Introduction
Our interest is in investigating conditions for the existence of solutions to general nonconvex minimization problems by developing a simple unified framework relying on asymptotic cones of functions and sets. The existence of solutions has been extensively studied starting with seminal work [11] showing that a quadratic function (bounded from below) attains its minimum on a polyhedral set. The result has been extended to the problem with a quasi-convex objective function in [14]. The work in [5] has established that a convex polynomial attains its solution on a region described by finitely many convex polynomial inequalities, which in turn generalized the result established in [20] for convex quadratic functions.
@@
The inherent difficulty in establishing the existence of solutions is due to directions in unbounded constraint set along which the function may decrease. A framework to address this developed in [1, 3, 2, 17], relying on concepts such as the asymptotic cone and asymptotic function to show existence of solutions. Subsequently, in [7], this condition on asymptotic directions has been extended to introduce retractive directions and prove existence of solutions via a nonempty level-set intersection approach. The work in [16] has developed the existence results for problems where the constraint sets are given by functional inequalities. More recently, a solution existence result has been provided in [12] for a general polynomial objective and a closed constrained set under a certain regularity condition. Sufficient conditions for existence of solutions to convex problems and convex-concave saddle point problems over convex sets can be found in [6]. Recently, work in [10] has used generalized asymptotic functions to obtain sufficient conditions for the existence of solutions to equilibrium problems in an infinite-dimensional space and, as a special result, provides a sufficient condition for the existence of solutions to a minimization problem.
Similar to the works [1, 3, 2, 16, 17] we make use of the asymptotic cones and asymptotic functions. Unlike the aforementioned work, we introduce a concept of asymptotically bounded decay of a function over a given constraint set (with respect to a coercive function) and use it to establish necessary and sufficient conditions for the existence of solutions. We introduce the cone of retractive directions for a set and a function and investigate their properties. While retractive directions for sets have been silently used to define asymptotically linear sets in [2], and also used in [7, 17] with a slightly different, but more general definition developed for a family of nested sets, the cone of such directions has not been studied. Using these cones and asymptotically bounded decay property we establish new necessary and sufficient conditions for existence of solutions. the properties of the cones of retractive directions for a set and a function. More specifically, the novel aspects of the work include:
(1) Defining a class of functions that have asymptotically bounded decay with respect to a coercive function, which can be viewed as generalization of the coercivity and super-coercivity [8, 13]. This class of functions includes proper closed convex functions and functions with Lipschitz continuous $p$-derivatives, with $p\geq 0$, such as general multivariate polynomial functions (Section 3.1).
(2) Introducing a cone of retractive directions of a set and a function, and exploring their properties. Interestingly, while nonempty, such cones need not be closed, as shown by an example (Section 3.2). We also explore the relationships of a cone of retractive directions of a function and that of a level set of the function. Surprisingly, while asymptotic cone of a level set is always contained in the asymptotic cone of the function, this is not true for the cone of retractive directions. Even more surprisingly, for two distinct lower-level sets of a function, there is no particular inclusion relation for the cones of their retractive directions, even when the function is convex. Finally, we provide a closed form expression for the asymptotic function of a nonconvex polynomial and the cone of retractive directions of a convex polynomial (Section 3.3).
(3) Establishing necessary and sufficient conditions for the existence of solutions based on asymptotically bounded decay property and a relation for cones of retractive directions of the objective function $f$ and the constraint set $X$ (Theorems 4.14.3, Theorem 4.5, Section 4). The paper  [17] is critical in showing the necessity of the conditions. We refine our results for general nonconvex problems to a more structured problems, such as convex, nonconvex with functional constraints and polynomial constraints (Sections 56), thus obtaining new results that are more general than those in [5, 16, 7].
@@
This paper is organized as follows: Section 2 provides the necessary background required for the subsequent development. Section 3 introduces the notion of asymptotically bounded decay of a function, the cones of retractive directions of a set and a function, and investigates their properties. Section 4 presents our main results and their proofs for the problems where the constraint set is generic and not specified via inequalities. Section 5 and Section 6 revisits our main results for the cases when the problem has more structure.
@@ 2 Notation and Terminology
We deal with the extended real-valued functions that take values in the set {$\mathbb{R}\cup\{+\infty\}$. For a function $f$, we use ${\rm dom}f$ to denote its effective domain, i.e., ${\rm dom}f=\{x\in\mathbb{R}^{n}\mid f(x)<+\infty\}$. The epigraph of a function $f$ is denoted by ${\rm epi}f$, i.e., ${\rm epi}f=\{(x,c)\in\mathbb{R}^{n}\times\mathbb{R}\mid f(x)\leq c\}$. For any $\gamma\in\mathbb{R}$, the lower-level set of a function $f$ is denoted by $L_{\gamma}(f)$, i.e., $L_{\gamma}(f)=\{x\in\mathbb{R}^{n}\mid f(x)\leq\gamma\}.$
Our focus is on proper and closed functions, which are the functions with the epigraph ${\rm epi}f$ that is nonempty and closed set.
The function $f:\mathbb{R}^{n}\rightarrow\mathbb{R}\cup\{+\infty\}~{}$is said to be coercive on a set $X$ if
$\lim_{\begin{subarray}{c}\|x\|\to+\infty\\ x\in X\end{subarray}}f(x)=+\infty.$
The notion of a $p$-supercoercive, with $p\geq 1$, function has been introduced in [8, 13], which states that: a proper function $f$ is $p$-supercoercive if $\liminf_{\|x\|\rightarrow\infty}\frac{f(x)}{\|x\|^{p}}>0.$
@@ 2.1 Asymptotic Cones
We turn our attention to unbounded sets whose behavior “at infinity” is captured by their asymptotic cones (see [2]), which are also referred to as horizon cones [19]. To define these cones, we start with the definition of a sequence converging in a direction.
Definition 2.1.
A sequence $\{x_{k}\}\subset\mathbb{R}^{n}$ is said to converge in the direction $d\in\mathbb{R}^{n}$ if there exists a scalar sequence $\{t_{k}\}\subset\mathbb{R}$ with $t_{k}\rightarrow+\infty$ such that $\lim_{k\rightarrow\infty}t_{k}^{-1}x_{k}=d.$
Definition 2.2 (Definition 3.3 of [19]).
@@
For an empty set $X$, we have $X_{\infty}=\{0\}$.
The set $X_{\infty}$ is always nonempty closed cone (Proposition 2.1.1 in [2]). Asymptotic cones are illustrated in Figure 2.1 and Figure 2.2. We conclude this section with a result for the asymptotic cone of the intersection of sets (Proposition 2.1.9 in [2], or 3.9 Proposition in [19]).
Proposition 2.1.
@@
$(\cap_{i\in\mathcal{I}}C_{i})_{\infty}\subseteq\cap_{i\in\mathcal{I}}(C_{i})_{\infty}.$
@@
$x$$y$
Figure 2.1: Set $X=\{(x,y)\in\mathbb{R}^{2}\mid x^{2}\leq|y|\}$ and its asymptotic cone $X_{\infty}=\{(0,y)\mid y\in\mathbb{R}\}$.
$x$$y$
Figure 2.2: Set $X=\{(x,y)\in\mathbb{R}^{2}\mid x^{2}\geq y\}$ and its asymptotic cone $X_{\infty}=\mathbb{R}^{2}$.
@@ 2.2 Asymptotic Functions
Applying the concept of the asymptotic cone to the epigraph of a proper function leads the notion of an asymptotic function, which is also referred to as a horizon function[19]. We elect to define it according to [2], as stated below.
Definition 2.3.
@@
A useful analytic representation of an asymptotic function $f_{\infty}$ was originally obtained in [1] and, also, given in Theorem 2.5.1 of [2].
Theorem 2.2.
@@
$f_{\infty}(d)=\liminf_{\begin{subarray}{c}d^{\prime}\rightarrow d\\ t\rightarrow+\infty\end{subarray}}\frac{f(td^{\prime})}{t},$ (1)
or equivalently,
@@
An asymptotic function has some basic properties inherent from its definition, namely that $f_{\infty}$ is closed and positively homogeneous function since its epigraph is the closed cone $({\rm epi}f)_{\infty}$.
Further, the value $f_{\infty}(0)$ is either finite or $f_{\infty}(0)=-\infty$. If $f_{\infty}(0)$ is finite, then it must be that $f_{\infty}(0)=0$ by the positive homogeneity property. As a consequence, for a proper function $f$ we have that
$0\in\{d\mid f_{\infty}(d)\leq 0\},$
implying that $\{d\mid f_{\infty}(d)\leq 0\}\neq\emptyset.$ The directions $d$ such that $f_{\infty}(d)\leq 0$ will be particularly important in our subsequent development.
To this end, we will term them as asymptotic directions of a function, and use these directions to define the asymptotic cone of a function, as follows.
Definition 2.4.
@@
An asymptotic direction of a function has been given in Definition 3.1.2 of [2], while the asymptotic cone of a function has been defined for a proper convex function in Definition 2.5.2 of [2]. Here, we adopt the same definition for any proper function.
We next provide a key result for the asymptotic cones of lower-level sets of a proper function (Proposition 2.5.3 of [2] and 3.23 Proposition of [19]).
Proposition 2.3.
For a proper function $f$ and $\alpha\in\mathbb{R}$, we have $\left(L_{\alpha}(f)\right)_{\infty}\subseteq L_{0}(f_{\infty})$ i.e.,
@@
Finally, we state an existence of solutions result that will also be used in the sequel.
Proposition 2.4.
Let $X\subseteq\mathbb{R}^{n}$ be a closed set, and let $f:\mathbb{R}^{n}\rightarrow\mathbb{R}\cup\{+\infty\}~{}$be a proper closed function with $X\cap{\rm dom}f\neq\emptyset$. If $f$ is coercive over $X$, i.e.,
@@
The result follows by applying Theorem 2.14 of [4] to the function $f+\delta_{X}$, where $\delta_{X}$ is the characteristic function of the set $X$.
3 Asymptotically Bounded Decay, Cones of Retractive Directions
Here, we introduce a notion of asymptotically bounded decay of a function (Subsection 3.1), and define the cones of retractive directions for sets and functions and study their properties (Subsections 3.2 and 3.3, respectively).
@@ 3.1 Asymptotically Bounded Decay
We introduce a condition that prohibits the decrease of $f$ to infinity at a certain rate, which can be viewed as a generalization of the coercivity property.
Definition 3.1.
@@
The condition in (3) prohibits the function $f$ from approaching $\infty$- ∞, along the points in the set $X$, faster than the function $g$. Note that, if $f$ is $p$-supercoercive on $X$, with $p\geq 1$, then the bounded decay condition in (3) is satisfied with $g(x)=\|x\|^{p}$. The class of functions that have this asymptotically bounded decay is rather rich, as illustrated in the following examples.
@@
Thus, if $\inf_{x\in X}f(x)>-\infty$, then $f$ satisfies Definition 3.1 with $g(x)=\|x\|$. ∎
Next, we show that any proper convex function satisfies Definition 3.1 with $g(x)=\|x\|$ due to the linear underestimation property of a convex function.
@@
Let $f$ be a proper convex function such that $X\cap{\rm dom}f\neq\emptyset$, and let $x_{0}$ be a point in the relative interior of ${\rm dom}f$. Then, by Theorem 23.4 of [18], the subdifferential set $\partial f(x_{0})$ is nonempty. Thus, by the convexity of $f$ we have for a subgradient $s_{0}$ of $f$ at the point $x_{0}$ and for all $x\in X$,
@@
Our next example shows that if a function $f$ satisfies $f_{\infty}(d)\geq 0$ for all nonzero $d\in\mathbb{R}^{n}$, then $f$ satisfies Definition 3.1 with $g(x)=\|x\|$.
@@
$\liminf_{\|x\|\to\infty\atop x\in X}\frac{f(x)}{\|x\|}=\liminf_{\|x\|\to\infty}\frac{f(x)+\delta_{X}(x)}{\|x\|}=\liminf_{\begin{subarray}{c}d=x\|x\|^{-1}\\ \|x\|\to\infty\end{subarray}}\frac{f(\|x\|d)+\delta_{X}(\|x\|d)}{\|x\|},$
where $\delta_{X}(x)=0$ when $x\in X$ and otherwise, $\delta_{X}(x)=+\infty$. Therefore,
$\liminf_{\|x\|\to\infty\atop x\in X}\frac{f(x)}{\|x\|}$ $\geq\liminf_{d^{\prime}\to d\atop t\to\infty}\frac{f(td^{\prime})+\delta_{X}(td^{\prime})}{t}$
$\geq f_{\infty}(d)+(\delta_{X})_{\infty}(d).$
Since ($\delta_{X})_{\infty}(d)=\delta_{X_{\infty}}(d)$ and $f_{\infty}(d)\geq 0$ for all nonzero $d\in X_{\infty}$, it follows that
$\liminf_{\|x\|\to\infty\atop x\in X}\frac{f(x)}{\|x\|}\geq 0,$
and $f$ exhibits asymptotically bounded decay on $X$ with respect to $\|x\|$. ∎
Finally, we show that a $p$-times differentiable function, with Lipschitz continuous $p$th derivatives, satisfies Definition 3.1 with $g(x)=\|x\|^{p+1}$.
@@
where $D^{p}f(x)$ denotes the $p$th derivative of $f$ at a point $x$. Then, by Equation (1.5) of [15], we have that
@@
where we used the fact that $\lim_{\|x\|\to\infty}\Phi_{x_{0},p}(x)/\|x\|^{p+1}=0$.
Hence, $f$ exhibits asymptotically bounded decay on $X$ with respect to $g(x)=\|x\|^{p+1}$. ∎
As a special case of Example 3.4, a function with Lipschitz continuous gradients exhibits asymptotically bounded decay with respect to the function $g(x)=\|x\|^{2}$.
Moreover, note that multivariate polynomials are a special class of functions that fall under Example 3.4. In particular, a multivariate polynomial of order $m$, with $m\geq 1$, has a constant $m$th derivative, say equal to $B$. Thus, the $(m-1)$st derivative is Lipshitz continuous with the constant $B$. By Example 3.4, a multivariate polynomial has asymptotically bounded decay with respect to the function $g(x)=\|x\|^{m}$.
3.2 Cone of Retractive Directions of Sets